OpenStudy (anonymous):

Working on Rolle's Theorem right now: C(x) = 10(1/x + (x/(x+3))) Find the derivative and solve for x in the derivative.

OpenStudy (anonymous):

@RadEn Can you help?

OpenStudy (noelgreco):

To have a question on Rolle's Theorem you need two points at which the function has the same value, then show that f'(x) has a value of zero at some intermediate point.

OpenStudy (anonymous):

Ok, I solved C(3) = C(6)

OpenStudy (noelgreco):

OK. Now calculate C'(x), and show that it has a zero between 3 and 5.

OpenStudy (anonymous):

That's the part I'm having trouble with. I keep messing up somehow because when I solve for it, I get imaginary numbers

OpenStudy (anonymous):

\[C'(x) = 10[-1/x^2 - x^2/(x+3)^2]\] That's the derivative I get

OpenStudy (noelgreco):

Is the function truly\[\frac{ 10 }{ \left( \frac{ 1 }{ x }+\frac{ x }{ x+3 } \right) }\]?

OpenStudy (anonymous):

No, it's the denominator times the numerator

OpenStudy (noelgreco):

Opps, OK. I'm calculating the derivative.

OpenStudy (noelgreco):

C(x) = 10(1/x + (x/(x+3)) You failed to apply the quotient rule to the second term in parentheses.The whole derovative should be: \[10(\frac{ -1 }{ x ^{2} }+\frac{ 3 }{ \left( x+3)^{2} \right) })\] should be

OpenStudy (anonymous):

Ok, I'm working it out again now

OpenStudy (anonymous):

Alright, I got how you did the derivative, but now I'm having trouble solving for x.

OpenStudy (anonymous):

I got to \[1/3 = x^2/(x+3)^2\]

OpenStudy (noelgreco):

Get a common denominator for the parenthetical term, and set the numerator( it's a quadratic) to zero. The Quadratic Formula will give you two zeros: one of them is between 3 and 6.

OpenStudy (anonymous):

Ok, thank you. I'll work it out

OpenStudy (noelgreco):

Let me know if you can't get ~4.098

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I'm not getting that

OpenStudy (anonymous):

Actually, nevermind, I got it :D

OpenStudy (anonymous):

Thanks! :)